OpenStudy (anonymous):

cos^2-sin^2= (sqrt 3/2) need help guys, please

OpenStudy (campbell_st):

well cos(2x) = cos^(x) - sin^2(x) so then \[\cos(2x) = \frac{\sqrt{3}}{2}\] so use an exact value triangle to find 2x, then halve it to find x... hope it helps

OpenStudy (campbell_st):

oops should read \[\cos(2x) = \cos^2(x) - \sin^2(x)\]